What volume is occupied by 0.324 mol of CO2 at 212.5 K and 819 mmHg?
Answer #1

Given:

P = 819 mm Hg

= (819/760) atm

= 1.0776 atm

n = 0.324 mol

T = 212.5 K

use:

P * V = n*R*T

1.0776 atm * V = 0.324 mol* 0.08206 atm.L/mol.K * 212.5 K

V = 5.2428 L

Answer: 5.24 L